OpenStudy (reverdeparadis):

|a| - 3/4 = -5/8 a.) 1/8 or -1/8 b.) 7/8 or -7/8 c.) 1 3/8 or -1 3/8 d.) no solution

OpenStudy (anonymous):

hey are you still there

OpenStudy (azureilai):

-5/8+3/4 -5/8+6/8 =1/8 a= +/-1/8

OpenStudy (anonymous):

|a| = 1/8 so a = 1/8 or -1/8

OpenStudy (reverdeparadis):

thank you! :)